OpenStudy (anonymous):

between what two consecutive integers does the square root of 445 lie?

OpenStudy (raden):

sqrt(445) = 21.09...... (use the calculator if you feel hard) so, the two consecutive integers of this, is 21 and 22

OpenStudy (yttrium):

\[21^2 = 441\] \[22^2 = 484\] Hence between 21 and 22.